a man can complete a piece of work in 9 days but with the help of his son, he can do it in 6 days.the taken by his son to do it alone is
Answer by nerdybill(7384)

a man can complete a piece of work in 9 days but with the help of his son, he can do it in 6 days.the taken by his son to do it alone is
.
Let x = time (days) it takes the son alone
then
6(1/9 + 1/x) = 1
multiplying both sides 9x:
6(x + 9) = 9x
6x + 54 = 9x
54 = 3x
18 days = x